# MSCA-Net: Multi-Scale Context Aggregation Network for Infrared Small Target Detection

### v1

@@ Abstract
Detecting infrared small targets in complex backgrounds remains a challenging task because of the low contrast and high noise levels inherent in infrared images. These factors often lead to the loss of crucial details during feature extraction. Moreover, existing detection methods have limitations in adequately integrating global and local information, which constrains the efficiency and accuracy of infrared small target detection. To address these challenges, this paper proposes a novel network architecture named MSCA-Net, which integrates three key components: Multi-Scale Enhanced Detection Attention mechanism(MSEDA), Positional Convolutional Block Attention Module (PCBAM), and Channel Aggregation Block (CAB). Specifically, MSEDA employs a multi-scale feature fusion attention mechanism to adaptively aggregate information across different scales, enriching feature representation. PCBAM captures the correlation between global and local features through a correlation matrix-based strategy, enabling deep feature interaction. Moreover, CAB redistributes input feature channels, facilitating the efficient transmission of beneficial features and further enhancing the model’s detection capability in complex backgrounds. The experimental results demonstrate that MSCA-Net achieves outstanding small target detection performance in complex backgrounds. Specifically, it attains mIoU scores of 78.43%, 94.56%, and 67.08% on the NUAA-SIRST, NUDT-SIRST, and IRTSD-1K datasets, respectively, underscoring its effectiveness and strong potential for real-world applications.
@@ I Introduction
Infrared remote sensing technology plays a critical role in various applications, including military early warning [1], disaster early warning [2], secure night surveillance [3], traffic monitoring [4], aerospace imaging [5] and agriculture [6]. In the military domain, it facilitates the rapid identification of enemy targets, enhances battlefield situational awareness, and enables the early detection of potential threats, thereby improving the accuracy and efficiency of strategic decision-making [7]. In disaster rescue operations, infrared remote sensing precisely locates trapped individuals, supporting swift search and rescue efforts, significantly enhancing rescue efficiency and reducing casualties [8]. In environmental monitoring, this technology is employed for meteorological observations, atmospheric pollution detection, early warning of forest fires, and other applications. This allows for the timely detection of potential hazards and provides accurate data support to relevant agencies, ensuring that disaster response measures are both prompt and effective [9]. In the field of agriculture, infrared remote sensing is equipped with unmanned aerial vehicles to acquire remote sensing data of rice and wheat, and combined with a neural network model, provides intelligent solutions for optimizing crop harvesting schemes and monitoring diseases and pests [10, 11]. Moreover, owing to its passive operation, robust target recognition capabilities, and all-weather functionality, infrared remote sensing also holds significant value in small target detection.
@@
However, despite its considerable application potential, infrared small target detection still faces numerous practical challenges. First, in long-range imaging, small targets typically occupy only a minute fraction of the image—with contrast levels below 0.15, signal-to-noise ratios under 1.5, and overall pixel coverage of less than 0.15% [12]—resulting in a lack of distinct visual features such as color and texture, which makes segmentation particularly challenging in complex backgrounds. Furthermore, under the interference of complex, dynamically changing backgrounds and noise from wave and cloud effects, these small targets are often prone to blurring or occlusion, further complicating detection. Consequently, effectively suppressing background noise, enhancing the model’s feature extraction capabilities, and achieving multi-scale detection of small targets remain challenging tasks.
To address these issues, we propose the MSCA-Net model and introduce a multi-scale attention mechanism (MSEDA) to increase the model’s capacity for multi-scale feature extraction while fostering deep interactions between local and global information. In addition, a Positional Convolutional Block Attention Module (PCBAM) [13] is incorporated to strengthen contextual information extraction. To further augment the model’s representational capability, a channel aggregation module (CAB) [14] is also introduced. Our goal is to enhance multi-scale feature extraction, facilitate the integration of local and global information, and improve the overall feature expression of the model, thereby tackling the challenge of precise target segmentation amid background noise and complex interference.
Our study demonstrates that the proposed MSCA-Net model significantly improves the segmentation accuracy for infrared small targets, providing valuable insights for the application of deep learning in the field of infrared imaging. Specifically, the main contributions of this work are as follows:
@@
(1) We propose an infrared small target detection network, MSCA-Net, which integrates multi-scale feature detection and contextual information interaction mechanisms. This network adaptively extracts critical features across different scales while facilitating cross-layer feature fusion, significantly improving target detection performance and segmentation accuracy in complex backgrounds.
@@
(2) We design a Multi-Scale Attention Module (MSEDA), which employs a multi-scale feature fusion attention mechanism to adaptively aggregate information across different scales. This mechanism enriches feature representation by highlighting key information, thereby enhancing the model’s adaptability to complex scenarios and improving detection accuracy.
@@
(3) We introduce the PCBAM module and the Channel Aggregation Block (CAB). PCBAM combines the Position Attention Module (PAM) and the cross-space channel interaction mechanism (CBAM) to achieve cross-layer feature fusion effectively, thereby enhancing the model’s focus on key regions. Moreover, CAB strengthens contextual information interaction through channel aggregation, facilitating efficient transmission of beneficial features and further improving the model’s target detection capability in complex backgrounds.
@@
The remainder of this paper is organized as follows. Section 2 reviews related previous work. Section 3 details the specifics of our detection network. Section 4 describes the experimental setup, results, and analysis. Finally, Section 5 concludes the paper.
@@ II Related work
We can broadly classify existing infrared small target detection methods into two paradigms: model-driven approaches centered on physical feature modeling and data-driven approaches using deep learning [15]. This section briefly reviews the application of traditional model-driven methods as well as current data-driven methods in the segmentation of infrared small targets, focusing particularly on the significant research progress achieved by the U-Net architecture and its variants in this field.
@@ II-A Model-driven approaches
Over the past several decades, researchers have proposed a variety of model-driven infrared small target detection (IRSTD) algorithms. These methods encompass techniques inspired by the human visual system (HVS) [16, 17], low-rank approximation approaches [18, 19, 20], top-hat filtering [21], and local contrast-based algorithms [22]. They typically rely on explicit mathematical models and assumptions, demonstrating excellent performance when there is a pronounced contrast between the target and its background. However, these traditional approaches exhibit certain limitations, especially when targets are embedded in complex backgrounds or when loud background noise is present. Specifically, when targets are interfered with by intricate backgrounds or high-contrast noise, the detection performance of conventional methods often deteriorates markedly, leading to a high rate of false alarms. This degradation arises because model-driven methods usually assume a distinct contrast difference between the target and the background—a condition that is frequently violated in real-world scenarios. Consequently, in dynamic and complex environments, particularly in practical applications such as military reconnaissance and security surveillance, these traditional methods often fail to achieve the required accuracy.
@@ II-B Data-driven approaches
To overcome the limitations of conventional model-driven approaches, researchers have shifted their focus toward data-driven deep learning methods. Compared with traditional techniques, these methods exhibit a distinct advantage in suppressing background noise and have achieved promising results in infrared small target detection.
@@
Researchers initially introduced convolutional neural networks (CNNs) [23] for infrared small target detection, significantly enhancing detection performance because of their powerful feature extraction capabilities. This breakthrough preliminarily validated the immense potential of data-driven models in this domain. However, the conventional CNN down-sampling process, while high-level features are extracted, also results in a gradual loss of spatial information, posing a substantial challenge for the precise localization of small targets. Moreover, target segmentation plays a pivotal role in infrared small target detection. Unlike simple target detection tasks, segmentation requires not only the identification of the target but also the accurate delineation of its boundaries, which is crucial for the effective recognition and tracking of infrared small targets. Owing to their lack of pixel-level prediction mechanisms, traditional CNNs models and standard classification networks are often inadequate in meeting the stringent requirements for precise contour extraction inherent in target segmentation tasks.
@@
To overcome these issues, researchers have adopted the encoder–decoder U-Net architecture to address the shortcomings of conventional CNNs in small target detection. Originally introduced by Ronneberger et al. [24] for medical image segmentation, U-Net’s distinctive encoder–decoder design excels in precise segmentation tasks. By merging high-level feature maps with low-level ones through skip connections, the architecture effectively restores the spatial details of targets. This capability enables U-Net to accurately delineate small target regions in infrared images and maintain robust performance against complex backgrounds. Consequently, U-Net has gradually become the predominant approach in infrared small target detection, and its numerous variants (e.g., 3D-U-Net [25] and Attention U-Net [26]) have further enhanced both detection and segmentation performance. For example, in 2019, Wang et al. [27] proposed MDvsFA-cGAN, an infrared small target segmentation framework based on a generative adversarial network (GAN) paradigm that balances the trade-off between missed detections and false alarms. In 2021, Dai et al. [28] designed an asymmetric contextual modulation (ACM) module specifically for infrared small target detection to supplement bottom-up modulation channels. In the same year, Dai et al. [29] introduced ALCNet, which reconstructs traditional local contrast measurement methods into an end-to-end network incorporating attention modulation. In 2022, Zhang et al. [30] developed the edge block aggregator network ISNet, which is aimed at enhancing edge information across multiple layers to improve the contrast between background and targets.
@@
Nevertheless, a standalone U-Net architecture still struggles with segmentation accuracy in infrared small target detection. To address this limitation, researchers have combined U-Net with advanced attention mechanisms, thereby further boosting detection performance and segmentation precision by enhancing the model’s focus on key regions. For example, Ren D et al. [31] proposed DNANet, a densely nested attention network based on the U-Net++ architecture that recovers clear images directly from blurred inputs via a novel connection topology. Similarly, Wu X et al. [32] refined U-Net to develop a “U-Net within U-Net” framework (UIU-Net), which embeds a smaller U-Net into a larger U-Net backbone and incorporates both channel and spatial attention for multi-level, multi-scale object representation learning. In addition, Sun et al. [33] designed the RDIAN network, which integrates receptive field and direction-guided attention mechanisms; it extracts features from various regions through convolutional layers with different receptive fields while using a multi-directional attention mechanism to enhance target information in low-level feature maps. Hou et al. [34] introduced ISTDU-Net, an infrared small target detection network that converts single-frame infrared images into pixel-wise target probability maps. To strengthen the representation of small target features, the network incorporates feature map groups during down-sampling; fully connected layers, added within skip connections, effectively suppress background interference, improving target-background contrast. Xu et al. [35] further contributed to SCTransNet, a single-branch real-time segmentation network that employs a transformer to align semantic information with CNNs’ features; this network maintains the fast inference of lightweight single-branch CNNs while achieving high precision.
In addition to U-Net-based architectures, other network models have also played pivotal roles in infrared small target detection. For example, Zhao et al. [36] proposed APLCNet, a detection network model based on attention and partially learnable convolution. By designing a projection global self-attention module (PGSAM) that leverages IMT projection saliency, the model effectively encodes long-range features to capture contextual relationships between targets and their backgrounds, thereby increasing detection accuracy. Ma et al. [37] introduced MDCENet, a multi-dimensional cross-enhancement network that employs a multi-dimensional complementary transformer (MDCT) module to fuse local convolutional features with global transformer features. They also developed residual omni dimensional convolutional blocks (RODCBs) and a global composite feature fusion (GCFF) module to capture subtle target details and integrate contextual information from encoding, decoding, and shallow layers. In addition, Cao et al. [38] presented the YOLO-TSL model, which improves detection accuracy by incorporating a ternary attention-based backbone, and reduces computational complexity via a Slim-Neck structure combined with a VoV-GSCSP module. This method further optimizes bounding box regression through an internal MPDIoU loss, achieving more efficient infrared target detection while enhancing overall performance.
@@ III Methodology
Figure 1 illustrates the overall architecture of the MSCA-Net model. The U-Net baseline forms the foundation of the model, which also incorporates the MSEDA, PCBAM, and CAB modules. Initially, the MSEDA module introduces a multi-scale attention mechanism to bolster feature extraction capabilities. Moreover, the integration of the PCBAM module allows for more precise capture of small target positional information in images, while also enhancing the extraction of contextual information to achieve deep interactions between global and local features. Finally, the application of the CAB module augments the model’s decoding ability by mapping feature maps through channel concatenation and 1×1 convolution layers, thereby facilitating accurate pixel-wise classification.
@@
III-A Multi-scale enhanced detection attention module

### v2

@@ Abstract
In complex environments, detecting tiny infrared targets has always been challenging because of the low contrast and high noise levels inherent in infrared images. These factors often lead to the loss of crucial details during feature extraction. Moreover, existing detection methods have limitations in adequately integrating global and local information, which constrains the efficiency and accuracy of infrared small target detection. To address these challenges, this paper proposes a network architecture named MSCA-Net, which integrates three key components: Multi-Scale Enhanced Dilated Attention mechanism (MSEDA), Positional Convolutional Block Attention Module (PCBAM), and Channel Aggregation Feature Fusion Block (CAB). Specifically, MSEDA employs a multi-scale feature fusion attention mechanism to adaptively aggregate information across different scales, enriching feature representation. PCBAM captures the correlation between global and local features through a correlation matrix-based strategy, enabling deep feature interaction. Moreover, CAB enhances the representation of critical features by assigning greater weights to them, integrating both low-level and high-level information, and thereby improving the model’s detection performance in complex backgrounds. The experimental results demonstrate that MSCA-Net achieves strong small target detection performance in complex backgrounds. Specifically, it attains mIoU scores of 78.43%, 94.56%, and 67.08% on the NUAA-SIRST, NUDT-SIRST, and IRTSD-1K datasets, respectively, underscoring its effectiveness and strong potential for real-world applications.
@@ I Introduction
In multi-field application scenarios, infrared small target detection(IRSTD) plays an irreplaceable role, including military early warning [1], disaster early warning [2], secure night surveillance [3], traffic monitoring [4], aerospace imaging [5] and agriculture [6]. In the military domain, it facilitates the rapid identification of enemy targets, enhances battlefield situational awareness, and enables the early detection of potential threats, thereby improving the accuracy and efficiency of strategic decision-making [7]. In disaster rescue operations, IRSTD precisely locates trapped individuals, supporting swift search and rescue efforts, significantly enhancing rescue efficiency and reducing casualties [8]. In environmental monitoring, this technology is employed for meteorological observations, atmospheric pollution detection, early warning of forest fires, and other applications. This allows for the timely detection of potential hazards and provides accurate data support to relevant agencies, ensuring that disaster response measures are both prompt and effective [9]. In the field of agriculture, IRSTD is equipped with unmanned aerial vehicles to acquire remote sensing data of rice and wheat, and combined with a neural network model, provides intelligent solutions for optimizing crop harvesting schemes and monitoring diseases and pests [10, 11]. Moreover, owing to its passive operation, robust target recognition capabilities, and all-weather functionality, IRSTD also holds significant value in small target detection.
@@
However, despite its considerable application potential, IRSTD still faces numerous practical challenges.In long-distance imaging application scenarios, weak and small targets often present pixel-level representations—specifically manifested as contrast characteristics weaker than 0.15, signal-to-noise ratio levels less than 1.5, and pixel proportions not exceeding 0.15% [12]—resulting in a lack of distinct visual features such as color and texture, which makes segmentation particularly challenging in complex backgrounds. Furthermore, under the interference of complex, dynamically changing backgrounds and noise from wave and cloud effects, these small targets are often prone to blurring or occlusion, further complicating detection. Consequently, while effectively reducing background noise, enhancing the model’s ability in feature extraction and achieving multi-scale detection of small targets remains a challenging task.
To solve these problems, we proposed the MSCA-Net model and designed the Multi-Scale Enhanced Dilated Attention mechanism (MSEDA). By constructing a bidirectional feature interaction channel, the adaptive enhancement of multi-scale features and the deep coupling of local-global semantics were achieved. Secondly, the Positional Convolutional Block Attention Module (PCBAM) [13] is introduced, and the spatial - channel - positional triple attention mechanism is adopted to enhance the feature responses of the key regions. Finally, this study embedded the Channel Aggregation Feature Fusion Block (CAB) [14]. By assigning greater weighting of the critical information channels, the fusion of key features is effectively enhanced. These three modules form a collaborative optimization system, solving the problem of feature discrimination under dynamic background interference and providing a theoretical solution for precise positioning and segmentation in highly dynamic and complex backgrounds.
Experiments show that the multi-scale semantic enhancement network constructed in this study demonstrates significant advantages in the segmentation accuracy and detection efficiency of infrared dim and small targets, providing theoretical support and practical guidance for the innovative application of deep learning technology in the field of infrared imaging. Specifically, the contributions of this research are reflected in the following three aspects:
@@
(1) This study constructs a deep learning architecture for infrared dim and small target detection—MSCA-Net, based on cross-scale feature collaboration and context-aware modeling. This network adaptively extracts critical features across different scales while facilitating cross-layer feature fusion, significantly improving target detection performance and segmentation accuracy in complex backgrounds.
@@
(2) We design a Multi-Scale Enhanced Dilated Attention Module (MSEDA), which employs a multi-scale feature fusion attention mechanism to adaptively aggregate information across different scales. This mechanism significantly enhances the robustness of the model in environmental perception in complex scenarios by dynamically strengthening the representation of key semantics, and ultimately achieves a breakthrough in pixel-level detection accuracy.
@@
(3) We introduce the PCBAM module and design the Channel Aggregation Feature Fusion Block(CAB). PCBAM combines the Position Attention Module(PAM) and the cross-space channel interaction mechanism(CBAM) to achieve cross-layer feature fusion effectively, thereby enhancing the model’s focus on key regions. Moreover, CAB facilitates contextual information interaction by aggregating and weighting critical channels, facilitating efficient transmission of beneficial features and further improving the model’s target detection capability in complex backgrounds.
@@
The remainder of this paper is organized as follows. Section 2 reviews related previous work. Section 3 details the specifics of our detection network. Section 4 constructing the experimental verification system and demonstrating the effectiveness of the proposed algorithm through the method of quantitative division. Finally, Section 5 concludes the paper.
@@ II Related work
We can broadly classify existing IRSTD methods into two paradigms: model-driven approaches centered on physical feature modeling and data-driven approaches using deep learning [15]. This section briefly reviews the application of traditional model-driven methods and current data-driven methods in the segmentation of infrared small targets, focusing particularly on the significant research progress achieved by the U-Net architecture and its variants in this field.
@@ II-A Model-driven approaches
Over the past several decades, researchers have proposed a variety of model-driven IRSTD algorithms. These methods encompass techniques inspired by the human visual system (HVS) [16, 17], low-rank approximation approaches [18, 19, 20], top-hat filtering [21], and local contrast-based algorithms [22]. They typically rely on explicit mathematical models and assumptions, demonstrating excellent performance when there is a pronounced contrast between the target and its background. However, these traditional approaches exhibit certain limitations, especially when targets are embedded in complex backgrounds or when loud background noise is present. Specifically, when targets are interfered with by intricate backgrounds or high-contrast noise, the detection performance of conventional methods often deteriorates markedly, leading to a high rate of false alarms. This degradation arises because model-driven methods usually assume a distinct contrast difference between the target and the background—a condition that is frequently violated in real-world scenarios. Consequently, in dynamic and complex environments, particularly in practical applications such as military reconnaissance and security surveillance, these traditional methods often fail to achieve the required accuracy.
@@ II-B Data-driven approaches
To overcome the limitations of conventional model-driven approaches, researchers have shifted their focus toward data-driven deep learning methods. Compared with traditional techniques, these methods exhibit a distinct advantage in suppressing background noise and have achieved promising results in IRSTD.
@@
Researchers initially introduced convolutional neural networks (CNNs) [23] for IRSTD, significantly enhancing detection performance because of their powerful feature extraction capabilities. This breakthrough preliminarily validated the immense potential of data-driven models in this domain. However, the conventional CNN down-sampling process, while high-level features are extracted, also results in a gradual loss of spatial information, posing a substantial challenge for the precise localization of small targets. Moreover, Unlike simple target detection tasks, segmentation requires not only the identification of the target but also the accurate delineation of its boundaries, which is crucial for the effective recognition and tracking of infrared small targets. Owing to their lack of pixel-level prediction mechanisms, traditional CNNs models and standard classification networks are often inadequate in meeting the stringent requirements for precise contour extraction inherent in target segmentation tasks.
@@
To overcome these issues, researchers have adopted the encoder–decoder U-Net architecture to address the shortcomings of conventional CNNs in small target detection. Originally introduced by Ronneberger et al. [24] for medical image segmentation, U-Net’s distinctive encoder–decoder design excels in precise segmentation tasks. By merging high-level feature maps with low-level ones through skip connections, the architecture effectively restores the spatial details of targets. This capability enables U-Net to accurately delineate small target regions in infrared images and maintain robust performance against complex backgrounds. Consequently, U-Net has gradually become the predominant approach in IRSTD, and its numerous variants (e.g., 3D-U-Net [25] and Attention U-Net [26]) have further enhanced both detection and segmentation performance. For example, in 2019, Wang et al. [27] constructed a generative adversarial network architecture called MDvsFA-cGAN, which is based on the GAN paradigm and aims to achieve accurate segmentation of small infrared targets. In 2021, Dai et al.[28] designed an asymmetric context modulation architecture (ACM), which enhanced the feature interaction integrity of the bottom-up modulation channel by establishing an inverse feature compensation path, effectively solving the multi-scale semantic mismatch problem in traditional methods. In the same year, Dai et al. [29] proposed ALCNet, which transforms the classic local contrast analysis method into a local attention-guided method. Next year, Zhang et al. [30] developed the edge block aggregator network ISNet, aiming to construct cross-level edge features and significantly reduce the context-target semantic gap.
Nevertheless, a standalone U-Net architecture still struggles with segmentation accuracy in IRSTD. To resolve this bottleneck, researchers have integrated the U-Net architecture with advanced attention mechanisms and introduced a cross-level feature collaboration strategy to enhance the model’s responsiveness to critical information regions, thereby further improving the accuracy of infrared small target recognition and segmentation. For example, Ren D et al. [31] constructed DNANet, a dense nested attention network based on the U-Net++ architecture, which can realize an end-to-end network from degraded space to clear image space through topological connections. Similarly, Wu X et al. [32] improved U-Net and constructed a recursive nested attention framework (UIU-Net), which recursively embeds the U-Net subnet into the backbone U-Net and combines channel and spatial attention for multi-level and multi-scale object representation learning.
@@
In addition, Sun et al. [33] designed the RDIAN network, which combines receptive field and direction-aware attention mechanisms; it achieves feature capture of multi-scale receptive domains through convolution kernel configuration, while using multi-directional attention mechanisms to enhance the target information in the shallow feature space. Hou et al. [34] introduced ISTDU-Net, an infrared small target detection network that converts single-frame infrared images into pixel-wise target probability maps. To strengthen the representation of small target features, the network incorporates feature map groups during down-sampling; fully connected layers, added within skip connections, effectively suppress background interference, improving target-background contrast. Xu et al. [35] proposed the SCTransNet network, which introduces the core module SSCA (Spatial-Embedded Single-Head Channel Cross-Attention) to enhance semantic interactions across different feature levels and effectively model long-range contextual dependencies in images, thereby significantly improving the performance of infrared small target detection and segmentation accuracy.
In addition to U-Net-based architectures, other network models have also played pivotal roles in infrared small target detection. For example, Liu et al. [36] constructed the multi-order U-shaped Mamba network, and the Multi-Order 2D-Selective Scan (MO-SS2D) module and Local-Guided 2D-Selective Scan (LG-SS2D) module were constructed. The collaborative improvement of detection accuracy and reasoning efficiency has been achieved through the dynamic scanning mechanism of hierarchical feature interaction and local context guidance. Ma et al. [37] proposed integrating discrete Wavelet Transform (DWT) into a convolutional neural Network (CNN), and designed a hybrid attention mechanism, namely small waveguide direction Transformer (WST), to fully enhance the spatial and cross-channel characteristics of wavelet transform, achieving the collaborative optimization of detection accuracy and anti-noise performance. In addition, Yue et al. [38] presented the YOLO-MST model, which improves detection accuracy by incorporating super-resolution technology and multi-scale feature extraction, achieving more efficient infrared target detection while enhancing overall performance.
@@ III Methodology
Figure 1 presents the holistic structure of the MSCA-Net framework. The proposed model is built upon the U-Net baseline. Initially, four residual blocks (RBs) [39] and max-pooling layers are employed to extract images features $I_{i}\in\mathbb{R}^{C_{i}\times\frac{H}{i}\times\frac{W}{i}}$, ($i=1,2,3,4$). The corresponding channel dimensions are $C_{1}=32$, $C_{2}=64$, $C_{3}=128$, and $C_{4}=256$.
Next, patch embedding is applied to each $I_{i}$ using convolutions with kernel sizes and strides of $P$, $P/2$, $P/4$, and $P/8$, respectively. The resulting features are fed into the MSEDA and PCBAM modules for full-level semantic feature blending. Subsequently, to obtain $D_{i}\in\mathbb{R}^{C_{i}\times\frac{H}{16}\times\frac{W}{16}}$, ($i=1,2,3,4$), the features are then restored to the original encoder size through a feature mapping (FM) module [35], which integrates bilinear interpolation, convolution, batch normalization, and ReLU activation. In parallel, a residual connection is introduced to effectively bridge the semantic gap between the encoder $I_{i}$ and decoder $D_{i}$ by fusing their respective features. These representations are fused via the CAB module, followed by a CBL operation (Convolution + Batch Normalization + LeakyReLU). A final convolutional layer and a Sigmoid activation function are used to produce the prediction maps $T_{i}$, ($i=1,2,3,4$). Bilinear interpolation is employed to restore the image size of the output.
Finally, the Binary Cross-Entropy (BCE) loss [32] is used to calculate the difference between the predicted outputs and the ground truth (GT). The detailed structures of MSEDA, PCBAM, and CAB are described in Sections 3.1, 3.2, and 3.3, respectively.
@@
III-A Multi-scale enhanced dilated attention module